# Inside Out: Externalizing Assumptions in Data Analysis as Validation Checks

### v1

@@ 5.1 Data Simulation
To ensure a reasonable likeness to data that might be used in such an analysis, we use characteristics of the observed New York City dataset to refine our simulations. AIC is used to determine the best distribution fit for each variable with the QQ-plot presented in Figure 4 to evaluate the fit. AIC suggests a negative binomial distribution for mortality, a beta distribution for PM10 (multiple by 100 to recover the original scale), and a Weibull distribution for temperature. To include the potential effect of outliers, we add a single outlier to the data for both the mortality and PM10 variables.
Figure 4: QQ-plot of the distribution fit for mortality, PM10, and temperature based on the fitted distribution from the original data. The fitted distribution is compared to the observed data to assess the distribution fit.
A logic regression is fitted using all variations of the checks in Table LABEL:tbl-checks to predict whether the PM10 coefficient is unexpected. Given the imbalance of outcome (expected vs. unexpected), inverse weights proportional to the number of observations in each outcome are applied during the logic regression fit. Figure 5 shows the optimal logic regression tree from the fitted model. Precision, recall, and independence score, along with their harmonic and arithmetic mean are presented in Table 3.
Figure 5: Logic regression model fitted to the fourteen checks and the outcome expectation (unexpected) as the response variable. The model suggests the relationship: (mortality-PM10 correlation < -0.03) AND (mortality-temperature correlation > -0.3 OR (there exist mortality outlier AND there exist PM10 outlier)) to predict the unexpected PM10 coefficient.
@@
Check 1: cor(m, PM10) < -0.03 0.551 0.894 1 0.763 0.815
Check 2: cor(m, tmp) > -0.35 0.281 0.556 1 0.472 0.612
Check 3: mortality outlier 0.302 0.775 1 0.536 0.692
Check 4: PM10 outlier 0.305 0.766 1 0.537 0.690
Logic regression: (check 1) AND ((check 2) OR (check 3 AND check 4)) 0.722 0.876 1 0.851 0.866
@@
  • mortality-PM10 correlation less than $-0.03$
  • mortality-temperature correlation greater than $-0.3$
@@
(cor(mortality, PM10) < -0.03) AND (cor(mortality, tmp) > -0.3 OR (mortality outlier AND PM10 outlier))
This generates a precision of 0.722 and a recall of 0.876 for predicting the unexpected PM10 coefficient. As shown in Figure 5, there is no single analysis check in the tree that predicts an unexpected outcome. rather at least three checks in the tree must be TRUE in order for the model to predict an unexpected outcome. Given the high independence of the checks (Table 3), this suggests that unexpected results are only likely after multiple anomalies are observed in the data.
@@ 5.2 Additional Cities
Salt Lake City -0.0004 T T –0.19 (T) –0.03 (T) F T
San Antonio -0.0021 T T –0.22 (T) –0.16 (T) F T
@@
It is clear from Table 4 that some of the estimated PM10 coefficients are outside the expected range of [0, 0.005]. Eight of the 40 cities had negative estimated coefficients and were therefore considered unexpected by our original criterion. (It is perhaps worth noting that none of the unexpected outcomes was in the positive direction.) Table 5 summarizes the prediction accuracy of the logic regression model for the 40 cities. The model correctly predicts the PM10 coefficient status of 32 cities, producing an accuracy rate of 80%. Among the 8 cities with an unexpected outcome, the model correctly identifies 7 (Atlanta, Austin, Memphis, Orlando, Salt Lake City, San Antonio, Tampa), giving a recall of 88% (only Miami was not properly classified by the logic regression model). These cities were flagged due to failures in both the mortality-PM10 correlation and mortality-temperature correlation analysis checks. Out of the 14 cities with positive predictions from the model, 7 cities (Cincinnati, Columbus, Dallas/Fort Worth, Pittsburgh, Riverside, San Bernardino, Santa Ana/Anaheim) were false positives, resulting in a precision of only 50%. This precision value is substantially lower than what was estimated by the simulation procedure (see Table 3) and suggests that the simulation does not adequately capture some features of the data generation process.
@@ 6 Discussion
In this paper we have developed an approach to using analysis validation checks to externalize the assumptions about the data and analysis tools made during the data analysis process. These checks can serve as a useful summary of the analyst’s thought process and can describe how characteristics of the data may lead to unexpected outcomes. Using logic regression, we can develop a graphical summary of the analysis validation checks as well as use the logic regression fitting process to choose the optimal set of checks. The logic regression model can also be used to develop summaries of the precision and recall of the collection of analysis validation checks in predicting the likelihood of an unexpected outcome. We demonstrated our method on an example relating daily mortality to outdoor air pollution data. The results from that example could be used to inform future analyses of air pollution and health data, perhaps in other cities or locations.
@@
An interesting connection can be drawn between our logic regression trees and a tool used in systems engineering known as a fault tree. A fault tree is used for conducting a structured risk assessment and has a long history in aviation, aerospace, and nuclear power applications [Vesely et al., 1981]. A fault tree is a graphical tool that describes the possible combinations of causes and effects that lead to an anomaly. At the top of the tree is a description of an anomaly. The subsequent branches of the tree below the top event indicate possible causes of the event immediately above it in the tree. The tree can then be built recursively until we reach a root cause that cannot be further investigated. Each level of the tree is connected together using logic gates such as AND and OR gates. The leaf nodes of the tree indicate the root causes that may lead to an anomaly. While the logic regression tress are not identical to fault trees, they share many properties, such as the tree-based structure and the indicator of root causes at the leaf nodes. Perhaps more critically, both serve as graphical summaries of the assumptions made in a problem and the specific violations of those assumptions that could lead to an unexpected result. While fault trees are often used to discover the root cause of an anomaly after it occurs, an important use case for fault trees is to develop a comprehensive understanding of a system before an anomaly occurs [Michael et al., 2002].
@@
Visualization methods are also valuable tools for assessing data assumptions and can potentially be formalized as analysis validation checks. For instance, plotting a variable’s distribution using a histogram, density plot, or bee swarm plot can reveal outliers or deviations from normality. These visualizations could be re-framed as analysis checks that fail when the data does not conform to the visual expectation. However, translating visualization results into binary checks remains an open challenge, requiring either manual verification or the development of automated methods to interpret visualization outputs. An existing example of visual test is the R package vdiffr [Henry et al., 2023] for graphic software unit testing. The package saves a template plot and compares it to the current plot to determine whether the unit tests passes or fails.
@@
Systematically generating realistic simulated data is a key component of our approach and is deserving of further consideration. In the PM10 example in Section 5, the inverse transform method was used to preserve the correlation structure among mortality, PM10, and temperature. However, the simulation process can become complex when additional restrictions are imposed or when a greater range of scenarios is desired. In such cases, techniques like the acceptance-reject method or permutation may be used to generate the data. The results in Section 5 on the broader NMMAPS dataset suggest that our simulation procedure may have been inadequate in reflecting the range of possible configurations that the data could take. In particular, our split-data approach, using New York City to guide the simulation and then applying the analysis checks to other cities, may not have been ideal. It may be worth exploring some recent work in data thinning [Neufeld et al., 2024] and data fission [Leiner et al., 2023] to create training datasets that are more representative of future data.
@@
Communicating the process of data analysis is a key element to providing transparency, improving reproducibility, and building trust with a range of audiences. The analysis validation checks described here provide a general way to encode the assumptions that an data analyst makes about the data and statistical tools applied to the data. These assumptions can be studied or challenged, depending on the specific analyst’s perspective, and can serve as a roadmap for diagnosing unexpected results.

### v2

@@ 5.1 Data Simulation
To ensure reasonable similarity to observed data, we use characteristics of the observed New York City dataset to refine our simulations. AIC is used to determine the best distribution fit for each variable with the QQ-plot presented in Figure 4 to evaluate the fit. AIC suggests a negative binomial distribution for mortality, a beta distribution for PM10 (multiplied by 100 to recover the original scale), and a Weibull distribution for temperature. To include the potential effect of outliers, we add a single outlier to the data for both the mortality and PM10 variables.
Figure 4: $Q$-plot of the goodness-of-fit of selected distributions used to generate simulated data for mortality, PM10, and temperature. Left: Mortality as a negative binomial distribution (size = 74, mean = 183), Middle: PM10 is modeled using a rescaled beta distribution with shape parameters 4.21 and 11.67, multiplied by 100 to match the observed range. Right: Temperature is modeled using a Weibull distribution (shape = 3.8, scale = 61)
A logic regression is fitted using all variations of the checks in Table LABEL:tbl-checks to predict whether the PM10 coefficient is unexpected. Given the outcome imbalance (expected vs. unexpected), inverse weights proportional to the number of observations in each outcome are applied during the logic regression fit. Figure 5 shows the optimal logic regression tree from the fitted model. Precision, recall, and independence score, along with their harmonic and arithmetic means are presented in Table 3.
Figure 5: Logic regression model fitted to the fourteen checks and the outcome expectation (unexpected) as the response variable. The model suggests the relationship: (mortality-PM10 correlation < 0.03-0.03) AND (mortality-temperature correlation > 0.3-0.3 OR (there exists mortality outlier AND there exists PM10 outlier)) to predict the unexpected PM10 coefficient.
@@
Check 1: cor(m, PM10) < – 0.03 0.551 0.894 1.000 0.763 0.815
Check 2: cor(m, tmp) > – 0.35 0.281 0.556 1.000 0.472 0.612
Check 3: mortality outlier 0.302 0.775 1.000 0.536 0.692
Check 4: PM10 outlier 0.305 0.766 1.000 0.537 0.690
Logic regression: (check 1) AND ((check 2) OR (check 3 AND check 4)) 0.876 0.722 0.812 0.798 0.803
@@
  • mortality-PM10 correlation less than 0.03-0.03
  • mortality-temperature correlation greater than 0.3-0.3
@@
(cor(mortality, PM10) < – 0.03) AND (cor(mortality, tmp) > – 0.3 OR (mortality outlier AND PM10 outlier))
This generates a precision of 0.876 and a recall of 0.722 for predicting the unexpected PM10 coefficient. As shown in Figure 5, there is no single analysis check in the tree that predicts an unexpected outcome. Rather, at least three checks in the tree must be TRUE in order for the model to predict an unexpected outcome. Given the high independence of the checks (Table 3), this suggests that unexpected results are only likely after multiple anomalies are observed in the data.
@@ 5.2 Additional Cities
Salt Lake City –0.0004 T T –0.19 (T) –0.03 (T) F T
San Antonio –0.0021 T T –0.22 (T) –0.16 (T) F T
@@
It is clear from Table 4 that some of the estimated PM10 coefficients are outside the expected range [0, 0.005]. Eight of the 40 cities had negative estimated coefficients and were therefore considered unexpected by our original criterion. (It is perhaps worth noting that none of the unexpected outcomes was in the positive direction.) Table 5 summarizes the prediction accuracy of the logic regression model for the 40 cities. The model correctly predicts the PM10 coefficient status of 32 cities, producing an accuracy rate of 80%. Among the 8 cities with an unexpected outcome, the model correctly identifies 7 (Atlanta, Austin, Memphis, Orlando, Salt Lake City, San Antonio, Tampa), giving a recall of 88% (only Miami were not properly classified by the logic regression model). These cities were flagged due to failures in both the mortality-PM10 correlation and mortality-temperature correlation analysis checks. Out of the 14 cities with positive predictions from the model, 7 cities (Cincinnati, Columbus, Dallas/Fort Worth, Pittsburgh, Riverside, San Bernardino, Santa Ana/Anaheim) were false positives, resulting in a precision of only 50%. This precision value is substantially lower than what was estimated by the simulation procedure (see Table 3) and suggests that the simulation does not adequately capture some features of the data generation process.
@@ 6 Discussion
In this paper, we have developed an approach to using analysis validation checks to externalize the assumptions about the data and analysis tools made during the data analysis process. These checks can serve as a useful summary of the analyst’s thought process and can describe how characteristics of the data may lead to unexpected outcomes. Using logic regression, we can develop a graphical summary of the analysis validation checks as well as use the logic regression fitting process to choose the optimal set of checks. The logic regression model can also be used to develop summaries of the precision and recall of the collection of analysis validation checks in predicting the likelihood of an unexpected outcome. We demonstrated our method on an example relating daily mortality to outdoor air pollution data. The results from that example could be used to inform future analyses of air pollution and health data, perhaps in other cities or locations.
@@
An interesting connection can be drawn between our logic regression trees and a tool used in systems engineering known as a fault tree. A fault tree is used for conducting a structured risk assessment and has a long history in aviation, aerospace, and nuclear power applications [Vesely et al., 1981]. A fault tree is a graphical tool that describes the possible combinations of causes and effects that lead to an anomaly. At the top of the tree is a description of an anomaly. The subsequent branches of the tree below the top event indicate possible causes of the event immediately above it in the tree. The tree can then be built recursively until we reach a root cause that cannot be further investigated. Each level of the tree is connected together using logic gates such as AND and OR gates. The leaf nodes of the tree indicate the root causes that may lead to an anomaly. While the logic regression trees are not identical to fault trees, they share many properties, such as the tree-based structure and the indicator of root causes at the leaf nodes. Perhaps more critically, both serve as graphical summaries of the assumptions made in a problem and the specific violations of those assumptions that could lead to an unexpected result. While fault trees are often used to discover the root cause of an anomaly after it occurs, an important use case for fault trees is to develop a comprehensive understanding of a system before an anomaly occurs [Michael et al., 2002].
@@
Visualization methods are also valuable tools for assessing data assumptions and can potentially be formalized as analysis validation checks. For instance, plotting a variable’s distribution using a histogram, density plot, or bee swarm plot can reveal outliers or deviations from normality. These visualizations could be re-framed as analysis checks, which fail when: the data does not conform to the visual expectation. However, translating visualization results into binary checks remains an open challenge, requiring either manual verification or the development of automated methods to interpret visualization outputs. An existing example of visual test is the R package vdiffr [Henry et al., 2023] for graphic software unit testing. The package saves a template plot and compares it to the current plot to determine whether the unit tests pass or fail.
@@
Systematically generating realistic simulated data is a key component of our approach and is deserving of further consideration. In the PM10 example in Section 5, the inverse-transform method was used to preserve the correlation structure among mortality, PM10, and temperature. However, the simulation process can become complex when additional restrictions are imposed or when a greater range of scenarios is desired. In such cases, techniques like the acceptance-reject method or permutation may be used to generate the data. The results in Section 5 on the broader NMMAPS dataset suggest that our simulation procedure may have been inadequate in reflecting the range of possible configurations that the data could take. In particular, our split-data approach, using New York City to guide the simulation and then applying the analysis checks to other cities, may not have been ideal. It may be worth exploring some recent work in data thinning [Neufeld et al., 2024], data fission [Leiner et al., 2023], or differential privacy methods [Dong et al., 2022] to create training datasets that are more representative of future data. The Gaussian copula used in the example doesn’t account for the tail dependence between the variables and other copula, e.g. t-copula or extreme value distribution copula, could also be explored to better capture the tail dependence.
@@
Communicating the process of data analysis is a key element to providing transparency, improving reproducibility, and building trust with a range of audiences. The analysis validation checks described here provide a general way to encode the assumptions that a data analyst makes about the data and statistical tools applied to the data. These assumptions can be studied or challenged, depending on the specific analyst’s perspective, and can serve as a roadmap for diagnosing unexpected results.
@@ 7 Acknowledgement
Supplementary material
The supplementary materials include a full script of the examples in the paper (index.R) and its output (index.html), the data used in the examples in Section 5 (data/), the package source (adtoolbox_0.1.0.tar.gz), and a README.md file containing the install instructions for running the scripts.